# Embodied Crowd Counting

### v1

@@ Abstract
Occlusion is one of the fundamental challenges in crowd counting. In the community, various data-driven approaches have been developed to address this issue, yet their effectiveness is limited. This is mainly because most existing crowd counting datasets on which the methods are trained are based on passive cameras, restricting their ability to fully sense the environment. Recently, embodied navigation methods have shown significant potential in precise object detection in interactive scenes. These methods incorporate active camera settings, holding promise in addressing the fundamental issues in crowd counting. However, most existing methods are designed for indoor navigation, showing unknown performance in analyzing complex object distribution in large scale scenes, such as crowds. Besides, most existing embodied navigation datasets are indoor scenes with limited scale and object quantity, preventing them from being introduced into dense crowd analysis. Based on this, a novel task, Embodied Crowd Counting (ECC), is proposed. We first build up an interactive simulator, Embodied Crowd Counting Dataset (ECCD), which enables large scale scenes and large object quantity. A prior probability distribution that approximates realistic crowd distribution is introduced to generate crowds. Then, a zero-shot navigation method (ZECC) is proposed. This method contains a MLLM driven coarse-to-fine navigation mechanism, enabling active Z-axis exploration, and a normal-line-based crowd distribution analysis method for fine counting. Experimental results against baselines show that the proposed method achieves the best trade-off between counting accuracy and navigation cost.
Figure 1: Comparisons between our dataset&model and others. (a) Comparison of environment areas between existing robot navigation datasets and ours. Our dataset features large-scale outdoor crowd scenes. (b) Comparison of lenses interactivity between existing crowd counting datasets and ours. Our dataset includes a 3D structural environment for each scenario, enabling interaction with agents. (c) Comparison of methods between traditional crowd counting models and ours. We propose using agent navigation to achieve precise detection in previously unexplored faraway areas.
[…]
@@ 1 Introduction
Crowd counting is critical for public safety and urban planning [21]. One main challenge in this field is occlusion. It can be categorized into two aspects: overlap and invisibility. Overlap refers to the high density of people stacked together, making it difficult to distinguish each individual from some viewpoints, while invisibility indicates that the current camera position is obstructed, such as being blocked by buildings, or far from the crowds so that the target cannot be detected clearly. To summarize, these situations are caused by a poor observation point. Existing methods try to solve these challenges from several perspectives, such as using multi-scale feature extraction [20], body part detection [26] or multi-camera fusion [49, 50]. Datasets are also expanded for enhancing method performance [51, 19, 36, 43]. In general, these methods or datasets either try to enhance learned model representation, or expand the camera sensing range by introducing multi-view settings. Yet, their passive camera settings do not fully solve the occlusion challenge in crowd counting, especially in cases where crowds exceed the sensing range, or no cameras are set to detect completely obscured crowds. These issues restrict passive-camera-based methods in practice.
@@
Recent development in Embodied AI brings perspectives to solve the occlusion challenge in crowd counting. It has been demonstrated to possess significant potential in enhancing scene exploration and object detection. Aspects such as Vision-Language-Navigation (VLN) have been directed towards equipping mobile robots with human-like perception abilities, resulting in remarkable performance in exploring scenes and detecting objects in open environments [9, 4, 52, 7]. The active camera settings in VLN is promising to solve the fundamental challenges in crowd counting, since it can optimize observation point to mitigate overlap and invisibility caused by poor camera position setting. Yet most VLN benchmarks [44, 40, 3, 1] are indoor environments with limited exploration space (e.g., no Z-axis action options) and relatively small object quantity. And the performance of such methods remains unknown for detecting crowds with a large quantity and complex distribution in large-scale scenes. This results in a significant gap between VLN and crowd counting, as in practice, crowds often appear in large spaces with varying distribution.
To address the issues mentioned above, we first define a novel task, Embodied Crowd Counting (ECC). This task integrates crowd counting into the Zero-shot Object-Goal Navigation task (ZSON) in VLN [28]. It is designed for large-scale scenes with a large object quantity. A testing environment, Embodied Crowd Counting Dataset (ECCD), is then developed for this task. Compared to existing crowd counting and VLN datasets, this test environment has the following features: 1) It supports 60 distinct and diverse large-scale outdoor scenes. Each scene covers an area up to 40,000 $m^{2}$, with target quantity up to 15,000. 2) It uses a prior distribution, Poisson Point Process [11], for crowd quantity distribution modeling, which approximates realistic crowd situations.
Based on this platform, we develop a method, Zero-shot Embodied Crowd Counting (ZECC), to conduct crowd counting in ECCD. Compared to existing crowd counting methods and VLN methods, ZECC has the following features: 1) It uses active camera setting to solve fundamental challenges in crowd counting, transferring the detection task to an exploration task; 2) It consists of a coarse-to-fine navigation mechanism which uses common sense of Multi-Modal-Large-Language-Model (MLLM) to analyze the surrounding environment and plan Z-axis exploration, leveraging advantages of 6 degrees of freedom (6-DoF) to avoid low-altitude obstacles and gain border sight, thus achieving efficient exploration; 3) It consists of a crowd distribution analysis mechanism which transfers the complex crowd distribution to simple surfaces. The normal lines of the surfaces are used to generate fine observation points, which are used to mitigate the influence of overlap and invisibility, enabling accurate crowd counting.
The contributions of this work can be summarized as follows:
@@
  • ZECC is proposed to use MLLM for Z-axis exploration, reducing costs while ensuring detection performance. By utilizing normal lines to calculate navigation points, this approach eliminates overlap and enhances visibility in crowds.
  • Experiments show that the proposed method outperforms the baseline methods in various testing environments.
2 Related Work
@@ 2.1 Crowd Counting
Crowd-counting algorithms have greatly benefited from large-scale, high-quality datasets like UCF-CC50 [18] and UCF-QNRF [19]. These foundational datasets have facilitated the creation of subsequent collections focused on dense crowd imagery, such as ShanghaiTech [51], JHU-CROWD++ [36], NWPU-Crowd [43]. However, the images in these datasets are generated from fixed cameras. In contrast, ECCD provides interactive capabilities while maintains diverse crowd distribution.
Methods like [51, 20, 24, 24, 5] leverage crowd distribution prior in an image, or using attention maps to learn dependency. Recent multi-modal approaches [10, 47, 29, 42] leverage vision-language models to transfer image-text knowledge to dense crowd prediction. While these models improve long range and overlapping small target detection ability, their performance are restricted in practical if overlap reaches extreme level. Recent efforts expand spatial coverage. These works include multi-view systems [32, 17] that use multi camera to capture images from larger scale scenes. Others like [14, 15] use recorded video to conduct crowd analysis. While these are important advances on solving the fundamental challenges in crowd counting against the single-fix-image-based methods, their camera setting is still predefined by the dataset and cannot be adjusted when inferring. This cannot ensure fully exploration in a large environment. ZECC allows active exploration, which is fundamentally different from existing methods.
@@ 2.2 Embodied Navigation
Many traditional robot navigation methods were developed using conventional navigation datasets like KITTI [12] and SUN RGB-D [38]. Beyond these foundational datasets, [44, 33] provide 3D indoor environments for navigation and interaction tasks. These datasets are limited with their small scale size and small object quantity. Recent datasets have been created towards outdoor navigation like AerialVLN [25], and CityNav [23]. However, they are designed for VLN tasks without consideration of object quantity. Compared to these datasets, ECCD simultaneously supports large scale outdoor scenes and large object quantity with diverse distribution.
Efficient exploration using mobile robots remains a crucial challenge in vision and robotics. [13, 16, 30, 38] have developed human-like cognitive maps, enabling autonomous path learning in unknown environments. Other approaches like [2, 6, 31] use reinforcement learning to develop exploration policies. Recently, [48] applied video-based visual-language models to plan sequential actions in VLN. And [27, 53] presented zero-shot models that utilizes natural language instructions to guide agents through environments without prior environment-specific training. However, the methods mentioned above are for indoors navigation without Z-axis moving ability. This restricts their ability in large scale outdoor scenes. Recently, methods like [25, 45] introduce 6-DoF in outdoor VLN. Yet they are designed for instruction following task, in which the movement is restricted by human language. Besides, they lack ability to analyze crowds with large quantity and complex distribution. In contrast, ZECC is the first self-motivated agent in 6-DoF and can handle crowds simultaneously.
Figure 2: The proposed framework for ZECC consists of several key steps. First, we utilize an exploration strategy based on FBE to create a global crowd distribution by generating a target point cloud. Next, we apply a Gaussian Mixture Model (GMM) clustering to identify the representative centers of the global target point cloud and determine the final navigation points based on potential navigation directions. Finally, we integrate detection results obtained from various detectors with depth observations, mapping these to the global point cloud. The number of filtered targets is then counted and recorded as the final result for the crowd.
@@ 3.1 Problem Definition
To ensure the interactive with environment for accurately crowd counting in vast outdoor environments, we propose an innovative task called Embodied Crowd Counting (ECC). In ECC, an agent is first deployed in an unknown environment. At time step $t$, the input of the agent is RGB-D observations ${O_{t}}$ along with its pose ${p_{t}}$. Based on this information, the agent predicts navigation point $p\subset\mathbb{R}^{3}$ to a drone at time step $t$ and the drone moves to $p$. During exploration, the agent is allowed to record observations that help with crowd counting. When the agent decides to stop, it outputs an integer that represents its crowd counting result. The counting error and travel distance are calculated to assess the agent’s performance. ECC can be considered analogous to the ZSON task, since they both require an agent to detect targets in an unexplored environment without additional assistance. However, there are several differences. The task setting differences are: 1) The object category $G$ in ZSON is limited to crowd in ECC. 2) The step limitation in ZSON is removed in ECC. 3) The evaluation criterion in ECC is changed to counting error and travel distance. These differences do not prevent the availability of most recent ZSON methods in ECC by conducting minor modifications. However, agents should face essential differences in ECC: 1) Z-axis is available, which is not considered by most recent ZSON methods. 2) Complex crowd distributions exist, including heavy occlusions, while ZSON methods only consider objects with small quantity. Current ZSON methods show unknown performance under such differences, and new methods should be considered in ECC. Note that ECC is different from the instruction following tasks in VLN such as [25, 45], since these tasks require natural language assistance to drive the agent.
@@ 3.2 Embodied Crowd Counting Dataset (ECCD)
We propose a new dataset called ECCD to support algorithm design and evaluation for the community. ECCD is developed using Unreal Engine 4. This platform enables programmable environment construction, allowing scene richness and scalability. The characteristics of ECCD are as follows:
Diversity. Since different environments have different layouts, crowd distributions, and quantities reflect different challenges, ECCD is designed to contain diverse scenes. This dataset contains 60 distinct environments. Also, ECCD has an area up to 40,000 $m^{2}$, reaches a height up to 50 $m$, and allows for the simultaneous presence of over 15,000 targets within a single scene. This is significantly different from existing robot navigation datasets, such as [33] which offers environments with an average navigable space of 1,000 $m^{2}$, and crowd counting datasets, such as [51] which contains 1,198 annotated images captured with static cameras and a total of 330,165 individuals.
Realism. The ECCD is designed to simulate large-scale outdoor crowd scenarios in reality, in order to ensure practical effectiveness of systems built on ECCD, as shown in Figure 3. Environments like a city, a stadium, and a parking lot are simulated. Additionally, the environments support a complex structure of buildings in real life, such as multiple floors and bleachers. These features allow ECCD to reflect challenges in reality.
Figure 3: ECCD is designed to realistically mimic building and crowd distribution. On the left are samples from ECCD, and on the right are the real scenes.
Crowd generation mechanism. To model real challenges in crowd counting, ECCD uses Poisson Point Process for crowd quantity distribution modeling [11]. In ECCD, Blocks are placed in the simulator to present potential areas where crowds may exist. Then, for each block $\mathcal{U}\subset\mathbb{R}^{2}$, the process is defined as:

### v2

@@ Abstract
Occlusion is one of the fundamental challenges in crowd counting. In the community, various data-driven approaches have been developed to address this issue, yet their effectiveness is limited. This is mainly because most existing crowd counting datasets on which the methods are trained are based on passive cameras, restricting their ability to fully sense the environment. Recently, embodied navigation methods have shown significant potential in precise object detection in interactive scenes. These methods incorporate active camera settings, holding promise in addressing the fundamental issues in crowd counting. However, most existing methods are designed for indoor navigation, showing unknown performance in analyzing complex object distribution in large-scale scenes, such as crowds. Besides, most existing embodied navigation datasets are indoor scenes with limited scale and object quantity, preventing them from being introduced into dense crowd analysis. Based on this, a novel task, Embodied Crowd Counting (ECC), is proposed to count the number of persons in a large-scale scene actively. We then build up an interactive simulator, the Embodied Crowd Counting Dataset (ECCD), which enables large-scale scenes and large object quantities. A prior probability distribution approximating a realistic crowd distribution is introduced to generate crowds. Then, a zero-shot navigation method (ZECC) is proposed as a baseline. This method contains an MLLM-driven coarse-to-fine navigation mechanism, enabling active Z-axis exploration, and a normal-line-based crowd distribution analysis method for fine counting. Experimental results show that the proposed method achieves the best trade-off between counting accuracy and navigation cost. Code can be found at https://github.com/longrunling/ECC?.
@@ 1 Introduction
Crowd counting is critical for public safety and urban planning kang2018beyond . One main challenge in this field is occlusion. It can be categorized into two aspects: overlap and invisibility. Overlap refers to the high density of people stacked together, making it difficult to distinguish each individual from some viewpoints, while invisibility indicates that the current camera position is obstructed, such as being blocked by buildings, or far from the crowds so that the target cannot be detected clearly. To summarize, these situations are caused by a poor observation point. Existing methods try to solve these challenges from several perspectives, such as using multi-scale feature extraction kang2018crowd , body part detection liu2019point , or multi-camera fusion zhang2019wide ; zhang2024multi . Datasets are also expanded for enhancing method performance zhang2016single ; idrees2018composition ; sindagi2020jhu ; wang2020nwpu . In general, these methods or datasets either try to enhance learned model representation or expand the camera sensing range by introducing multi-view settings. Yet, their passive camera settings do not fully solve the occlusion challenge in crowd counting, especially in cases where crowds exceed the sensing range, or no cameras are set to detect completely obscured crowds. These issues restrict passive-camera-based methods in practice.
Figure 1: (a) Comparison between ECCD and existing embodied navigation datasets. ECCD features large-scale outdoor crowd scenes. (b) Comparison between ECCD and crowd counting datasets. ECCD enables interactive ability. (c) Comparison between ZECC and existing crowd counting methods. ZECC is an agentic framework with automatic camera adjusting ability.
@@
Recent development in Embodied AI brings a new perspective to address the occlusion challenge in crowd counting. It has been demonstrated to possess significant potential in enhancing scene exploration and object detection. Aspects such as Vision-Language-Navigation (VLN) have been directed towards equipping mobile robots with human-like perception abilities, resulting in remarkable performance in exploring scenes and detecting objects in open environments gadre2023cows ; chen2023typefly ; zhao2023agent ; choi2024find . The active camera settings in VLN are promising to solve the fundamental challenges in crowd counting, since it can optimize observation points to mitigate overlap and invisibility caused by fixed camera position settings. Yet most VLN benchmarks wang2024embodiedscan ; wald2019rio ; chang2017matterport3d ; brazil2023omni3d are indoor environments with limited exploration space (e.g., no Z-axis action options) and relatively small object quantity. And the performance of such methods remains unknown for detecting crowds with a large quantity and complex distribution in large-scale scenes. This results in a significant gap between VLN and crowd counting, as in practice, crowds often appear in large spaces with varying distribution.
To address the issues, we first define a novel task, Embodied Crowd Counting (ECC), which is shown in Figure 1. The task is defined as counting the total number of people present in a large outdoor scene using a drone. Given the absence of an existing dataset, we have created a new dataset called the Embodied Crowd Counting Dataset (ECCD) specifically for this task. This dataset includes 60 unique and diverse large-scale outdoor scenes. Each scene spans an area of up to 40,000 $m^{2}$ and has a target crowd size of up to 15,000 individuals. To ensure realism in our dataset, we employ a Poisson Point Process ge2009marked for modeling the distribution of crowd sizes, which effectively simulates real-world crowd scenarios. The comparision of ECCD and related datasets are shown in Table 1.
Table 1: Comparison between ECCD and other related datasets. ECCD combines features from both crowd counting datasets and embodied navigation datasets.
Dataset Active camera Place Max target quantity / sample Dynamic target w/o Instruction Task
NWPU-Crowd wang2020nwpu $\times$ - 20,033 $\times$ - $C$
SenseCrowd li2022video $\times$ - 296 $\checkmark$ - $C$
DukeMTMC ristani2016performance $\times$ - 2,834 $\checkmark$ - $C$
R2R anderson2018vision $\checkmark$ Indoor (Ground) - $\times$ $\times$ VLN
HM3DSem ramakrishnan2021habitat $\checkmark$ Indoor (Ground) - $\times$ $\checkmark$ ObjNav
AerialVLN gao2024aerial $\checkmark$ City(Aerial) - $\times$ $\times$ VLN
CityNav lee2024citynav $\checkmark$ City(Aerial) - $\times$ $\times$ VLN
Openfly gao2025openfly $\checkmark$ City(Aerial) - $\times$ $\times$ VLN
ECCD $\checkmark$ City(Aerial) 15,488 $\times$ $\checkmark$ ECC
In this study, we introduce a baseline method, Zero-shot Embodied Crowd Counting (ZECC) aiming at counting individuals present in environments populated by crowds. Given the ability of interacting dynamically with surroundings using foundation models, modern argentic methods achieve even better performance against training-based methods while maintaining generalization ability. gadre2023cows ; kuang2024openfmnav ; gao2024aerial ; long2024instructnav . Inspired by this, we aim to build a zero-shot baseline that can generate to diverse scenes leveraging foundation models, using an argentic paradigm without a vast amount of training data. The primary challenge lies in choosing suitable navigation points to find a balance between efficient exploration and effective detection. To address this, our approach consists of two main components: the Active Top-down Exploration (ATE) method and the Normal-line Based Navigation (NLBN) system. The ATE method serves as an efficient exploration strategy that employs a coarse-to-fine navigation approach. It utilizes the common sense capabilities of Multi-Modal Large Language Models (MLLM) to assess the environment, allowing for effective planning of vertical movement (Z-axis exploration). This takes advantage of the six degrees of freedom (6-DoF), which helps to avoid obstacles at lower altitudes and provides a broader view of the surroundings, thereby facilitating better exploration. Following the rough estimation of crowd distribution produced by ATE, we propose the NLBN to create precise navigation points for improved crowd detection. By using the normal lines of surfaces, we can establish detailed observation points that achieve a balanced trade-off between exploration efficiency and detection performance. This method addresses challenges related to overlapping individuals and visibility issues, ultimately leading to more accurate crowd counting. Experimental results show that the method is effective due to its interactive ability. The contributions of this work can be summarized as follows:
@@
  • We propose a baseline method, ZECC, using MLLM for Z-axis exploration, reducing costs while ensuring detection performance. By utilizing normal lines to calculate navigation points, this approach eliminates occlusion and enhances visibility in crowds.
2 Related works
@@ 2.1 Crowd Counting
Crowd-counting algorithms have greatly benefited from large-scale, high-quality datasets like UCF-CC50 idrees2013multi and UCF-QNRF idrees2018composition . These foundational datasets have facilitated the creation of subsequent collections focused on dense crowd imagery, such as ShanghaiTech zhang2016single , JHU-CROWD++ sindagi2020jhu , NWPU-Crowd wang2020nwpu . However, the images in these datasets are generated from fixed cameras. In contrast, ECCD provides interactive capabilities while maintaining diverse crowd distribution.
Methods like zhang2016single ; kang2018crowd ; liu2023point ; chen2024improving leverage crowd distribution prior in an image, or use attention maps to learn dependency. Recent multi-modal approaches gao2024clip ; yao2024cpt ; meng2024multi ; wang2024multi leverage vision-language models to transfer image-text knowledge to dense crowd prediction. While these models improve long-range and overlapping small target detection ability, their performance is restricted if the overlap reaches an extreme level. Recent efforts expand spatial coverage by multi-view systems qiu20223d ; hu2024mvd that use multiple cameras to capture images from large-scale scenes. Others like han2022dr ; han2023change use recorded video to conduct crowd analysis. Although these advances represent significant progress in addressing the basic challenges of crowd counting compared to traditional methods that rely on fixed images, the camera settings remain predetermined by the dataset. This restriction means that the settings cannot be modified during the inference process, limiting the ability to fully examine larger environments. ZECC allows active exploration, which is fundamentally different from existing methods.
@@ 2.2 Embodied Navigation
Many traditional robot navigation methods were developed using conventional navigation datasets like KITTI geiger2012we and SUN RGB-D song2015robot . Beyond these foundational datasets, wang2024embodiedscan ; ramakrishnan2021habitat provide 3D indoor environments for navigation and interaction tasks. These datasets are limited with their small scale size and small object quantity. Recent datasets have been created towards outdoor navigation liu2023aerialvln ; lee2024citynav ; gao2025openfly ; liu2025aerialvg However, they are designed for VLN tasks without consideration of object quantity. Compared to these datasets, ECCD simultaneously supports large-scale outdoor scenes and large object quantity with diverse distribution.
Efficient exploration using mobile robots remains a crucial challenge in vision and robotics. gupta2017cognitive ; henry2012rgb ; newcombe2011kinectfusion ; song2015robot ; gao2024aerial have developed human-like cognitive maps, enabling autonomous path learning in unknown environments. Other approaches like burda2018exploration ; chen2019learning ; parisi2021interesting use reinforcement learning to develop exploration policies. Recently, zhang2024navid applied video-based visual-language models to plan sequential actions in VLN. And long2024instructnav ; zhou2024navgpt presented zero-shot models that utilize natural language instructions to guide agents through environments without prior environment-specific training. However, the methods mentioned above are for indoor navigation without Z-axis moving ability. This restricts their ability in large-scale outdoor scenes. Recently, methods like liu2023aerialvln ; wang2024towards introduce 6-DoF in outdoor VLN. Yet they are designed for instruction following tasks, in which the movement is restricted by human language. Besides, they lack the ability to analyze crowds with a large quantity and complex distribution. In contrast, ZECC is the first self-motivated agent in 6-DoF that can handle crowds.
@@ 3.1 Problem Definition
To ensure the interactive with environment for accurately crowd counting in vast outdoor environments, we propose an innovative task called Embodied Crowd Counting (ECC). In ECC, an agent is first deployed in an unknown environment. At time step $t$, the input of the agent is RGB-D observations $O_{t}$} along with its pose $p_{t}$}. Based on this information, the agent predicts navigation point $p\subset\mathbb{R}^{3}$ to a drone at time step $t$ and the drone moves to $p$. During exploration, the agent is allowed to record observations that help with crowd counting. When the agent decides to stop, it outputs an integer that represents its crowd counting result. The counting error and travel distance are calculated to assess the agent’s performance. ECC can be considered analogous to the Zero-Shot Object-Goal Navigation (ZSON) task majumdar2022zson , since they both require an agent to detect targets in an unexplored environment without additional assistance. However, ECC faces unique challenges: 1) The Z-axis is available, which is not considered by most ZSON methods. 2) Complex crowd distributions exist, including heavy occlusions, while ZSON methods only consider fewer objects. Current ZSON methods show unknown performance under such differences, and new methods should be considered in ECC. Note that ECC is different from the instruction following tasks in VLN such as liu2023aerialvln ; wang2024towards , since these tasks require natural language assistance to drive the agent.
@@ 3.2 Embodied Crowd Counting Dataset (ECCD)
We propose a new dataset called ECCD to support algorithm design and evaluation for the community, developed using Unreal Engine 4. This platform enables programmable environment construction, allowing scene richness and scalability. The characteristics of ECCD are as follows:
Diversity. Since different environments have different layouts, crowd distributions, and quantities reflect different challenges, ECCD is designed to contain diverse scenes. This dataset contains 60 distinct environments. Also, ECCD has an area up to 40,000 $m^{2}$, reaches a height up to 50 $m$, and allows for the simultaneous presence of over 15,000 targets within a single scene. This is significantly different from existing robot navigation datasets, such as ramakrishnan2021habitat , which offers environments with an average navigable space of 1,000 $m^{2}$, and crowd counting datasets, such as zhang2016single , which contains 1,198 annotated images captured with static cameras and a total of 330,165 individuals.
Realism. The ECCD is designed to simulate large-scale outdoor crowd scenarios in reality, in order to ensure practical effectiveness of systems built on ECCD, as shown in Figure 2 (a). Environments like a city, a stadium, and a parking lot are simulated. Additionally, the environments support a complex structure of buildings in real life, such as multiple floors and bleachers. These features allow ECCD to reflect challenges in reality.
Figure 2: (a) ECCD is designed to mimic building and crowd distribution realistically. On the left are samples from ECCD, and on the right are the real scenes. (b) Illustration of the potential navigation vectors, normal lines, and FBE view vectors. Zoom in for better visualization.
Crowd generation mechanism. To model real distribution in crowd counting, ECCD uses Poisson Point Process for crowd quantity distribution modeling ge2009marked . In ECCD, Blocks are placed in potential areas where crowds may exist. Then, for each block $\mathcal{U}\subset\mathbb{R}^{2}$, the process is defined as:
@@ 3.3 Zero-shot Embodied Crowd Counting (ZECC)
Figure 3: The proposed framework. First, ATE is proposed to estimate the global crowd distribution efficiently. Then, NLBN is proposed to generate fine observation points, alleviating crowd overlap. The final result is generated by aggregating all fine detections.